Liquids A and B form an ideal solution. Calculate the total equilibrium vapour pressure and the mole fraction composition of the vapour in equilibrium with a solution of A and B containing 0.4753 mole fraction of B. P
1
Expert's answer
2021-08-12T06:43:49-0400

Raoult's law states that the vapor pressure of a solvent above a solution is equal to the vapor pressure of the pure solvent at the same temperature scaled by the mole fraction of the solvent present:



Total equilibrium vapour pressure = 1 - 0.4753=0.5247

Mole fraction = 0.4753 × 100 =47.53%